OpenStudy (anonymous):

It says to graph each linear equation how would i do these three : x+y=1 and x-y=-2 and x-2y=6

OpenStudy (anonymous):

use intercepts: set each variable equal to zero to find to others.

OpenStudy (anonymous):

I keep geting it wrong

OpenStudy (anonymous):

but first put them in to y=mx+b form

OpenStudy (anonymous):

I get -3+1/3

OpenStudy (anonymous):

How do put the numbers into the formula?

OpenStudy (anonymous):

you take the y out of the eqation sO all u have so far is Y=the u transfer the # so u get Y=X+1 (y=Mx+b )

OpenStudy (anonymous):

So how would i do mines, can you plz use mines as an example. I still cat get it correct..:(

OpenStudy (anonymous):

just keep it when u move the # so then it cahges to a negitive

OpenStudy (anonymous):

so X-Y=-2 will be -y=x-2 or y=-x+2 :)